On April 15, 2021, fire damaged the office and warehouse of Vaughn Corporation. The only accounting record saved was the general ledger, from which the balance sheet data below was prepared.

VAUGHN CORPORATION
MARCH 31, 2021

Cash

$21,640

Accounts receivable

41,310

Inventory, December 31, 2020

77,490

Land

34,680

Buildings

112,620

Accumulated depreciation

$44,180

Equipment

3,860

Accounts payable

22,809

Other accrued expenses

15,076

Common stock

103,100

Retained earnings

50,970

Sales revenue

134,520

Purchases

50,970

Miscellaneous expense

28,085

$370,655

$370,655


The following data and information have been gathered.

1. The fiscal year of the corporation ends on December 31.
2. An examination of the April bank statement and canceled checks revealed that checks written during the period April 1–15 totaled $12,890: $5,595 paid to accounts payable as of March 31, $3,438 for April merchandise shipments, and $4,219 paid for other expenses. Deposits during the same period amounted to $13,859, which consisted of receipts on account from customers with the exception of a $910 refund from a vendor for merchandise returned in April.
3. Correspondence with suppliers revealed unrecorded obligations at April 15 of $17,083 for April merchandise shipments, including $2,305 for shipments in transit (f.o.b. destination) on that date.
4. Customers acknowledged indebtedness of $47,140 at April 15, 2021. It was also estimated that customers owed another $8,080 that will never be acknowledged or recovered. Of the acknowledged indebtedness, $633 will probably be uncollectible.
5. The companies insuring the inventory agreed that the corporation’s fire-loss claim should be based on the assumption that the overall gross profit rate for the past 2 years was in effect during the current year. The corporation’s audited financial statements disclosed this information:

Year Ended
December 31

2020

2019

Net sales $487,600 $420,210
Net purchases 258,650 224,280
Beginning inventory 53,000 72,510
Ending inventory 77,490 53,000
6. Inventory with a cost of $7,640 was salvaged and sold for $3,520. The balance of the inventory was a total loss.


Compute the amount of inventory fire loss. (Round ratios for computational purposes to 2 decimal places, e.g 78.52% and final answer to 0 decimal places, e.g. 28,987.)
Answer #1

Inventory loss calculation

Opening inventory as on 1 Jan 2021 = $77,490

Purchases up to 31 Mar 2021 = $50,970

Less: Salvage goods = ($7,640)

Cost of sales up to 31 March 2021 = ($70,825) Calculated per below (A)

Inventory bought b/w 1-15 April 20121 = $2,528

(per info point 2, April shipments $3,438 Less refund $910 for April return)

Loss on account of salvage goods = $4,120

(Salvage of 7,640 less sold for $3,520)

Unrecorded obligations per info point 3 = $14,778

($17,083 minus $2,305)

Sales per Accounts receivable during 1-15 April less 47.35% margin as calculated below (A)

(Per info point 4 Balance as on 15 April $47,140 plus $8,080 less balance on 31 Mar 2021 $ 41,310 = $13,910 * (1-47.35%) = $7,324

Per info point 4 uncollectible debt = $633   

Hence, the total inventory loss due to fire = $79,378

Workings: (A) Calculation of % of gross profit based on 2019 & 2020 audited financial results

Net sales for both the years (just by adding two numbers given) = $907,810 (a)

Less: Cost of sales (Purchases plus beginning inventory less closing inventory) = $478,230 (b)

Gross profit = $429,860 (c=a-b)

% of gross profit = c/a = 47.35%